OpenStudy (anonymous):

1kg is equal to 100 g 10000 dg 1000 mg 1000 cg

OpenStudy (anonymous):

@midhun.madhu1987 @sammixboo

OpenStudy (anonymous):

10000 dg (1dg = 1/10g; 1g = 1/1000kg)

OpenStudy (anonymous):

dose that help

OpenStudy (anonymous):

Thanks. It does. ^^

OpenStudy (anonymous):

your welcome